A baseball has a mass of 0.150 kg, and it is pitched towards home plate at a speed of 56.0 m/s. Assume that the ball is pitched horizontally and that the ball is hit horizontally back in the opposite direction. If the bat exerts an average force of 8800 N for 2.20 ms, what is the final speed of the ball in m/s?
Answer #1

Solution)

We know,

Final momentum, Pf= mvf

Initial momentum, Pi= mvi=0.150*(-56)= -8.4 kg-m/s

Applying law of conservation of momentum,

Delta p= pf-pi

Pf= F*t + pi

Pf= (8800*0.0022) + (-8.4)

Pf= 10.96 kg-m/s

Now

Final speed, vf= Pf/m= 10.96/0.15= 73.06 m/s

=========